Diagnostic Supplies has expected sales of 176,400 units per year, carrying costs of $2 per unit, and an ordering cost of $4 per order. a. What is the economic ordering quantity? b-1. What is the average inventory? b-2. What is the total carrying cost? Assume an additional 70 units of inventory will be required as safety stock. c-1. What will the new average inventory be? c-2. What will the new total carrying cost be?
Answer #1

Assume Economic Order Quantity =Q

Annual Sales =176400

Number of Orders=176400/Q

Total Ordering Cost=(176400/Q)*$4

Average Inventory Level=Q/2

Total Carrying Cost=(Q/2)*$2

Total Cost =(Total Ordering Cost)+(Total Carrying Cost)

Total Cost will be minimum , when:

Total Ordering Cost=Total Carrying Cost

(176400/Q)*4=(Q/2)*2

Q^2=4*176400=705600

Q=705600^(1/2)=840

Economic Order Quantity

840

Average Inventory =Q/2=840/2=420

Average Inventory

420

Total Carrying Cost=(Q/2)*$2=(840/2)*$2=$840

Total Carrying Cost

$840

New Average Inventory =420+70=490

Total Carrying Cost =$840+70*$2=$980